OpenStudy (anonymous):

What is the simplified form of 1 over x-cubed all over 3 over x squared ? Explain please @SolomonZelman

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \frac{ ~~~\frac{1}{x^3}~~~~}{ \frac{3}{x^2}} }\) right ?

OpenStudy (anonymous):

Correct

OpenStudy (help!!!!):

its just like dividing fractions

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \frac{ ~~~\frac{1}{x^3}~~~~}{ \frac{3}{x^2}}=\frac{1}{x^3} \div\frac{3}{x^2} }\)

OpenStudy (anonymous):

Good what is next

OpenStudy (solomonzelman):

Then use \(\huge\color{blue}{ \frac{A}{B} \div\frac{C}{D}=\frac{A}{B} \times\frac{D}{C} }\)

OpenStudy (anonymous):

Will I multiply across?

OpenStudy (solomonzelman):

What exactly do you mean?

OpenStudy (solomonzelman):

you will multiply times x²/3. Because dividing by 3/x² is same as multiplying by x²/3

OpenStudy (anonymous):

Will A be mutiply by D

OpenStudy (solomonzelman):

Yes, \(\LARGE\color{blue}{ \frac{A \times D}{B \times C} }\)

OpenStudy (anonymous):

So, the answer is 3/x^5

OpenStudy (solomonzelman):

(1 × x²) ------ x³ × 3

OpenStudy (solomonzelman):

x² / 3x³

OpenStudy (solomonzelman):

1/3x

OpenStudy (anonymous):

Thanks